# Optimistic Estimation of Convergence in Markov Chains with the Average-Mixing Time

### v2

@@ Abstract
Keywords— Ergodic Markov chain; average mixing time; countable state space; $\beta$-mixing.
@@ 1   Introduction
The total variation distance variant of the mixing time $t_{\mathsf{mix}}$ of a Markov chain $X_{1},\dots,X_{n}$ on a discrete state space $\mathcal{X}$, governed by a transition operator $P$, quantifies the time to convergence to its stationary distribution $\pi$. Two observations on the chain separated in time by about $t_{\mathsf{mix}}$ will be approximately independent and sampled from $\pi$. Estimates on the mixing time yield exponential concentration inequalities (Paulin, 2015), and in turn, generalization bounds for stable machine learning algorithms trained with Markovian data (Mohri and Rostamizadeh, 2010; Agarwal and Duchi, 2012; Oliveira et al., 2022), regret bounds for reinforcement learning algorithms (Ortner, 2020), and convergence diagnostics for Markov Chain Monte Carlo methods, which are widely used in mathematical physics (Metropolis et al., 1953; Hastings, 1970), computer science (Sinclair, 2012) and statistics (Chib and Greenberg, 1995).
@@
for estimation down to constant multiplicative error (Hsu et al., 2019, Theorem 3.1, Theorem 3.2), where $\left|\mathcal{X}\right|$ is the size of the space and $\pi_{\star}$ is the minimum stationary probability (refer to Section 1.4). The sample complexity lower bound in (1) is prohibitively large for vast state spaces, and even vacuous in the infinite case.
In this paper, we propose to address some of the above listed concerns about the worst-case $t_{\mathsf{mix}}$, by considering the average-mixing time $t_{\mathsf{mix}}^{\sharp}$ which recently came under the spotlight in (Díaz et al., 2024; Münch and Salez, 2023). Specifically, we will confirm that $t_{\mathsf{mix}}^{\sharp}$ is a more optimistic measure of convergence, in the sense where some chains can average-mix arbitrarily faster than they mix, even on a two-point space. Furthermore, we initiate the program of estimating $t_{\mathsf{mix}}^{\sharp}$ from a single trajectory and we demonstrate that it is statistically less costly than the problem pertaining to $t_{\mathsf{mix}}$. In fact, contrary to the worst-case problem, our analysis carries to a large subset of countable-state Markov chains. One should rightfully expect $t_{\mathsf{mix}}^{\sharp}$ to be less informative than $t_{\mathsf{mix}}$ since it quantifies a weaker notion of mixing. However, we confirm that knowledge of $t_{\mathsf{mix}}^{\sharp}$ remains of practical value in the sense where it can be used in known powerful decoupling techniques (Yu, 1994) and appear as a natural parameter controlling the large deviation rate for additive functionals. Along the way, we explore the connections between the average-mixing time and $\beta$-mixing, argued by Vidyasagar (2013); Mohri and Rostamizadeh (2010) to be “just the right” assumption for the analysis of machine learning algorithms with weakly-dependent data.
@@ 1.1   Related work
Average variants of the mixing time have recently made their apparition in the literature. For instance, Berestycki et al. (2018) analyzed random walks on the giant component of the Erdös–Rényi random graph $\mathcal{G}(n,p)$ with $p=\Theta(1/n)$, and showed that the mixing time when starting randomly from any vertex is the order $\mathcal{O}(\log n)$ while it was known to be of the order $\mathcal{O}(\log^{2}n)$ for the mixing time from a worst starting point. Following up, Díaz et al. (2024) extended the approach on a larger collection of graphs with small bottlenecks, and show applicability of their approach for the Newman–Watts small-world model and supercritically percolated expander graphs. In (Lovász and Kannan, 1999), the authors introduce the notion of a conductance averaged over subsets of states with various sizes, and demonstrate that it can lead to tighter convergence bounds for counting and volume computation problems. Finally, the average-mixing time $t^{\sharp}_{\mathsf{mix}}$, considered in (Münch and Salez, 2023) and the object of our study, is a measure of convergence of a Markov chain defined with respect to the chain being started from its stationary distribution (refer to Section 2), instead of the worst-case state.
In particular, we stress that the notion of average-mixing treated in this paper is different from the one discussed in Peres and Sousi (2015); Anderson et al. (2021), defined by taking an average over two successive steps on the chain.
@@ 1.2   Main contributions
We direct the reader to Section 1.4 for the notation.
@@ 1.2.1 Highlight the significance of the average-mixing time
Observing that the average-mixing time $t^{\sharp}_{\mathsf{mix}}$ recently analyzed by Münch and Salez (2023) corresponds to the stationary $\beta$-mixing time of the Markov chain,
$t^{\sharp}_{\mathsf{mix}}(\xi)=\operatorname*{arg\,min}_{s\in\mathbb{N}}\left\{\beta(s)\leq\xi\right\},$
we briefly demonstrate how the average-mixing time, together with an a priori known order of convergence, can be used to quantitatively capture the convergence rate of a Markov chain in lieu of the worst-case mixing time. Specifically, we show (Lemma 2.3) that the average-mixing time $t^{\sharp}_{\mathsf{mix}}(\delta)$ naturally appears in the sample complexity for controlling the deviation probability from zero of a centered function $f\colon\mathcal{X}\to[-1,1]$. Namely, for $\frac{1}{n}\sum_{t=1}^{n}f(X_{t})\leq\varepsilon$ to hold with probability at least $1-\delta$, we show that under sub-exponential $\beta$-mixing of order $b$, it is sufficient to have a trajectory of
$n\geq\frac{\left(e/b\right)^{1/b}}{b\varepsilon^{2}}t^{\sharp}_{\mathsf{mix}}\left(\delta/e\right)\operatorname{polylog}\left(\frac{1}{\delta},\frac{1}{\varepsilon}\right),$
where $\operatorname{polylog}$ is a poly-logarithmic function of its parameters and $e$ is the Euler number. Similarly, in the polynomial $\beta$-mixing setting of order $b$, we show that for some natural constant $C\in\mathbb{R}_{+}$,
@@
is sufficient.
1.2.2 Estimation of $\beta$-mixing coefficients
For a general time-discrete, time-homogeneous Markov chain over a countable state space, we establish a parametric rate of convergence for its individual $\beta$-coefficients from a single stationary trajectory of observations $X_{1},\dots,X_{n}$. Namely, for $s\in\mathbb{N}$, we construct an estimator $\widehat{\beta}(s)$ for $\beta(s)$, and we show (refer to Theorem 3.1) that the absolute mean deviation can be controlled as follows,
$\begin{split}\mathbb{E}_{\pi}\left|\widehat{\beta}(s)-\beta(s)\right|\leq 4\sqrt{\frac{\mathcal{B}_{p}^{(s)}\mathcal{J}_{p}^{(s)}}{\lfloor(n-1)/s\rfloor}},\end{split}$
with
$\mathcal{B}_{p}^{(s)}\leq\left\|\beta^{(s)}\right\|_{1/p}^{1/p},\qquad\mathcal{J}_{p}^{(s)}=\left\|Q^{(s)}\right\|_{1-1/p}^{1-1/p},$
and where for $x,x^{\prime}\in\mathcal{X}$, $Q^{(s)}(x,x^{\prime})=\pi(x)P^{s}(x,x^{\prime})$. Here, $\mathcal{B}_{p}^{(s)}$ is a mixing parameter roughly controlled from above by the $\ell_{1/p}$ quasi-norm of the sequence of $\beta$-mixing coefficients of the $s$-skipped Markov chain, and $\mathcal{J}_{p}^{(s)}$ is an entropic quantity pertaining to the $s$-skipped Markov chain. Crucially, the parameter $p\in[1,\infty)$ is free and will allow us to balance the two factors in the upper bound.
@@ 1.2.3 Estimation of the average-mixing time
In the uniformly ergodic setting, for $\xi\in(0,1)$ and $\delta\in(0,1)$, we show (Theorem 3.4) that a single path length of
$n\geq\frac{t_{\mathsf{mix}}}{\xi^{2}}\inf_{p\geq 1}\left\{p\mathcal{J}_{p,3\xi/4}\right\}\operatorname{polylog}\left(\frac{1}{\xi},\frac{1}{\delta},t_{\mathsf{mix}}\right),$
where $t_{\mathsf{mix}}$ is the mixing time of $P$ and
$\mathcal{J}_{p,\xi}\doteq\sup\left\{\left\|Q^{(s)}\right\|_{1-1/p}^{1-1/p}\colon s\in\mathbb{N},s\leq t^{\sharp}_{\mathsf{mix}}(\xi)\right\},$
is sufficient in order to obtain an estimate $\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)$ of the average-mixing time $t^{\sharp}_{\mathsf{mix}}(\xi)$ which satisfies
$\begin{split}\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)\in\left[t_{\mathsf{mix}}^{\sharp}(3\xi/2),t_{\mathsf{mix}}^{\sharp}(\xi/2)\right]\end{split}$
with probability at least $1-\delta$. Furthermore, for finite spaces, setting $\xi=1/4$ for simplicity, the parameter $p$ can be chosen so as to obtain an upper bound (refer to Corollary 4.2) of
$n\geq Ct_{\mathsf{mix}}\log\max\left\{\left|\mathcal{X}\right|,\frac{t_{\mathsf{mix}}}{\delta}\right\},$
where $C\in\mathbb{R}_{+}$ is a universal constant. One can thus demonstrably estimate with high probability the average-mixing time sub-linearly in the number of states, answering a question posed by Luca Zanetti and John Sylvester. This discovery stands in stark contrast with the worst-case mixing time estimation problem, or the relaxation time estimation problem, for which the sample complexity necessarily depends inversely on the minimum stationary probability (Hsu et al., 2019; Wolfer and Kontorovich, 2024).
1.3   Outline
Section 2 introduces the average-mixing time, illustrates its significance and analyzes some of its properties. In Section 2.2, under geometric ergodicity and time-reversibility assumptions, we establish an upper bound of $t^{\sharp}_{\mathsf{mix}}$ in terms of the relaxation time of the chain, which remains non-vacuous over countable state spaces and under suitable conditions on the stationary distribution. In Section 2.3 we recall some important connection between the average-mixing time and $\beta$-mixing properties of the chain.
Section 3 is dedicated to the problem of inferring the $\beta$-mixing coefficients and average-mixing time from a single trajectory of observations. In Section 3.1, we estimate the $\beta$-mixing coefficients of the chain; first individually and then as a sequence, with respect to the uniform norm. From Section 3.3, we specialize our analysis to the uniformly ergodic setting, and in Section 3.4 we deduce an upper bound for the problem of estimating the average-mixing time from a single trajectory.
In Section 4, we demonstrate that our contribution holds implications at various scales of state spaces and provide more explicit rates under additional structural assumptions on the underlying chain.
Technical proofs have been deferred to Section 5 for readability.
Finally, Section 6 introduces a class of Markov chain which illustrates some of our results.
@@ 1.4   Notation and setting
$\ell_{p}\doteq\left\{s\in\mathbb{R}^{\mathcal{X}}\colon\left\|s\right\|_{p}<\infty\right\}.$
The set of all probability distributions over $\mathcal{X}$, written $\mathcal{P}(\mathcal{X})$, corresponds to non-negative functions summing to unity. A time-discrete, time-homogeneous Markov chain
$X_{1},X_{2},X_{3}\dots$
on $\mathcal{X}$ is defined by a transition operator represented by an infinite stochastic matrix $P$
$P(x,x^{\prime})=\mathbb{P}\left(X_{t+1}=x^{\prime}|X_{t}=x\right).$
We write $\mathcal{W}(\mathcal{X})$ for the set of all infinite stochastic matrices over $\mathcal{X}$. For a finite trajectory of length $n\in\mathbb{N}$, started with an initial distribution $\mu\in\mathcal{P}(\mathcal{X})$ and evolving according to $P\in\mathcal{W}(\mathcal{X})$, the probability of observing $(x_{1},x_{2},\dots,x_{n})\in\mathcal{X}^{n}$ can be factored as
@@
By abuse of notation, matrix multiplication will retain its meaning in the countable setting. Namely, for $\mu\in\mathcal{P}(\mathcal{X})$ and $P\in\mathcal{W}(\mathcal{X})$, we will write,
@@
Extensions of notation to $Pf^{\intercal}$ with $f\in\ell_{1}$, or for $P^{t}$ are similarly defined. We say that $P$ is ergodic when $P$ is irreducible, aperiodic and positive-recurrent (Levin et al., 2009). In that case, there exists a unique stationary distribution $\pi$ verifying $\pi P=\pi$, and we can define the distance to stationarity
@@
We write $\mathcal{W}^{\star}(\mathcal{X})$ the set of uniformly ergodic Markov chains. We recall that the (worst-case) mixing time is defined as
$\begin{split}t_{\mathsf{mix}}(\xi)\doteq\operatorname*{arg\,min}_{t\in\mathbb{N}}\left\{d(t)\leq\xi\right\}.\end{split}$
It is furthermore customary to define
@@
Elements related directly to the notion of average-mixing time will be introduced in the next section for clarity of the exposition.

### v4

@@ Abstract
Keywords— Ergodic Markov chain; average-mixing time; countable state space; $\beta$-mixing.
@@ 1  Introduction
The total variation distance variant of the mixing time $t_{\mathsf{mix}}$ of a Markov chain $X_{1},\dots,X_{n}$ on a discrete state space $\mathcal{X}$, governed by a transition operator $P$, quantifies the time to convergence to its stationary distribution $\pi$. Two observations on the chain separated in time by about $t_{\mathsf{mix}}$ will be approximately independent and sampled from $\pi$. Estimates on the mixing time yield exponential concentration inequalities (Chung et al., 2012; Paulin, 2015), and in turn, generalization bounds for stable machine learning algorithms trained with Markovian data (Mohri and Rostamizadeh, 2010; Agarwal and Duchi, 2012; Oliveira et al., 2022), regret bounds for reinforcement learning algorithms (Ortner, 2020), and convergence diagnostics for Markov Chain Monte Carlo methods, which are widely used in mathematical physics (Metropolis et al., 1953; Hastings, 1970), computer science (Sinclair, 2012) and statistics (Chib and Greenberg, 1995).
@@
for estimation down to constant multiplicative error (Hsu et al., 2019, Theorem 3.1, Theorem 3.2), where $\left|\mathcal{X}\right|$ is the size of the space and $\pi_{\star}$ is the minimum stationary probability (refer to Section 1.3). The sample complexity lower bound in (1) is prohibitively large when $\pi_{\star}\propto\exp(-\left|\mathcal{X}\right|)$ for vast state spaces, and even vacuous in the infinite case.
In this paper, we propose to address some of the above listed concerns about the worst-case $t_{\mathsf{mix}}$, by considering the average-mixing time $t_{\mathsf{mix}}^{\sharp}$ which recently came under the spotlight in Díaz et al. (2024); Münch and Salez (2023). Specifically, we will first confirm that $t_{\mathsf{mix}}^{\sharp}$ is a more optimistic measure of convergence, in the sense where some chains can average-mix arbitrarily faster than they mix, even on a two-point space. We then initiate the program of estimating $t_{\mathsf{mix}}^{\sharp}$ from a single trajectory of observations and we demonstrate that the problem is statistically less costly than the one pertaining to $t_{\mathsf{mix}}$. In fact, contrary to the worst-case problem, our analysis carries to a large subset of countable-state Markov chains. One should rightfully expect $t_{\mathsf{mix}}^{\sharp}$ to be less informative than $t_{\mathsf{mix}}$ since it measures a weaker notion of mixing. However, we confirm that knowledge of $t_{\mathsf{mix}}^{\sharp}$ remains of practical value in the sense where it can be used in known powerful decoupling techniques (Yu, 1994) and appears as a natural parameter controlling the large deviation rate for additive functionals. Along the way, we explore the connections between the average-mixing time and $\beta$-mixing, argued by Vidyasagar (2013); Mohri and Rostamizadeh (2010) to be “just the right” assumption for the analysis of machine learning algorithms with weakly-dependent data.
@@ 1.1  Related work
Concentration inequalities for Markov chains typically involve a parameter, such as a spectral gap or the mixing time, which accounts for dependencies within the sample (Lezaud, 2001; A León and Perron, 2004; Dedecker and Gouëzel, 2015; Havet et al., 2020). Estimates on this parameter yield corresponding deviation bounds. However, this parameter is often unknown and its definition is restricted to geometric ergodic settings.
1.1.1 Estimation of mixing time in Markov chains
The program of estimating the (worst-case) mixing time of a Markov chain from a single trajectory of observations was initiated by Hsu et al. (2015). In the time-reversible setting, they obtained the first upper and lower bounds on the sample complexity of estimating the absolute spectral gap of a chain in multiplicative error, and constructed fully empirical confidence intervals for the quantity. The analysis in the reversible setting was later complemented in Hsu et al. (2019), introducing an amplification method to obtain a generally unimprovable upper bound on the problem. In the more challenging non-reversible setting, the absolute spectral gap is no longer a good proxy for the mixing time. In Wolfer and Kontorovich (2024), the authors propose to estimate Paulin’s pseudo-spectral gap (Paulin, 2015) instead, and devise a method which enjoys the same sample complexity as in the reversible setting, modulo logarithmic factors. In parallel, Wolfer (2024) put forward a contraction approach, using a generalization of Dobrushin’s coefficient, instead of spectral methods to construct fully empirical confidence intervals for the mixing time.
1.1.2 Average-mixing time
Average variants of the mixing time have recently made their apparition in the literature. For instance, Berestycki et al. (2018) analyzed random walks on the giant component of the Erdös–Rényi random graph $\mathcal{G}(n,p)$ with $p=\Theta(1/n)$, and showed that the mixing time when starting randomly from any vertex is the order $\mathcal{O}(\log n)$ while it was known to be of the order $\mathcal{O}(\log^{2}n)$ for the mixing time from a worst starting point. Following up, Díaz et al. (2024) extended the approach on a larger collection of graphs with small bottlenecks, and show applicability of their approach for the Newman–Watts small-world model and supercritically percolated expander graphs. In (Lovász and Kannan, 1999), the authors introduce the notion of a conductance averaged over subsets of states with various sizes, and demonstrate that it can lead to tighter convergence bounds for counting and volume computation problems. Finally, the average-mixing time $t^{\sharp}_{\mathsf{mix}}$, considered by Münch and Salez (2023) and the object of our study, is a measure of convergence of a Markov chain defined with respect to the chain being started from its stationary distribution (refer to Section 2), instead of the worst-case state.
In particular, we emphasize that the notion of average-mixing discussed in this paper differs from that in Peres and Sousi (2015); Hermon and Peres (2017, 2018); Anderson et al. (2021). In these references, averaging is performed by taking an average over two successive steps of the chain to address periodicity issues. In contrast, our analysis involves averaging over the starting state, similar to the definition provided by Münch and Salez (2023).
@@ 1.2  Main contributions
We only introduce the setup and minimal notation in this section. We direct the reader to Section 1.3 for the comprehensive set of definitions used in this paper. Let $X_{1},\dots,X_{n}$ be a time-homogeneous time-discrete $\pi$-stationary ergodic Markov chain with transition operator $P$ over a countable state space $\mathcal{X}$. Assume that we are interested in approximating the stationary mean of a real function $f$ by the sample mean computed over the trajectory $X_{1},\dots,X_{n}$. More specifically, we seek a bound on the deviation $\left|f-\mathbb{E}_{\pi}f\right|$. To achieve a bound in the Markovian setting similar to those for iid data, the required sample size is typically multiplied by the worst-case mixing time (Chung et al., 2012; Paulin, 2015). The underlying idea is that one must allow the chain sufficient time to reach its steady state before regarding the observation as approximately independent and drawn from the stationary distribution. However, after the chain has reached steady state for the first time, waiting for an additional mixing time addresses the fact that we may have ended up in the worst possible state. While this approach is conservative, it may be unnecessarily cautious.
@@ 1.2.1 Highlight the significance of the average-mixing time
Once the chain has reached steady state, since the state we reach should be distributed according to the stationary distribution, it might be preferable to consider the next time to stationarity in a more probabilistic way. With this in mind, the question arises whether we can replace the worst-case mixing time with the following more intuitive average version,
$t^{\sharp}_{\mathsf{mix}}(\xi)=\operatorname*{arg\,min}_{t\in\mathbb{N}}\left\{\sum_{x\in\mathcal{X}}\pi(x)\left\|e_{x}P^{t}-\pi\right\|_{\mathsf{TV}}\leq\xi\right\},$
recently analyzed by Münch and Salez (2023). Under this criterion, convergence is achieved faster. In fact convergence can occur arbitrarily faster, for denumerable spaces—in which the worst-case mixing may not even be defined—and even for binary state spaces. We illustrate how together with an a priori known order of convergence, $t^{\sharp}_{\mathsf{mix}}$ can be used to summarize the convergence rate of a Markov chain in lieu of its worst-case mixing time. Specifically, we show (refer to Lemma 2.1) that the average-mixing time $t^{\sharp}_{\mathsf{mix}}(\delta)$ can be naturally plugged into the sample complexity for controlling the deviation probability from zero of a $1$-bounded centered function $f$. Indeed, for $\frac{1}{n}\sum_{t=1}^{n}f(X_{t})\leq\varepsilon$ to hold with probability at least $1-\delta$, we show that under sub-exponential $\beta$-mixing of order $b$ (Definition 2.1), it is sufficient to have a trajectory of observations of length
$n\geq\frac{\left(e/b\right)^{1/b}}{b\varepsilon^{2}}t^{\sharp}_{\mathsf{mix}}\left(\delta\right)\operatorname{polylog}\left(\frac{1}{\delta},\frac{1}{\varepsilon}\right),$
where $\operatorname{polylog}$ is a poly-logarithmic function of its parameters and $e$ is the Euler number. It is instructive to note that $b=1$ corresponds to the assumption that the chain is uniformly ergodic, and that the bound continues to hold for $b<1$, in which case the rate $b$ affects only the constant in the above expression. Similarly, in the polynomial $\beta$-mixing setting of order $b$ (Definition 2.1), we show that for some natural constant $C\in\mathbb{R}_{+}$,
@@
is sufficient. The trajectory lengths listed above both assume a stationary start. However, this assumption can be relaxed (refer to Section A).
@@ 1.2.2 Estimation of the average-mixing time
In the case where $\mathcal{X}$ is a finite space, for $\xi\in(0,1)$ and $\delta\in(0,1)$, our chief technical contribution is to show (refer to Corollary 4.1) that a trajectory $X_{1},X_{2},\dots,X_{n}$ of observations of length
$n\geq\frac{t_{\mathsf{mix}}}{\xi^{2}}\left|\mathcal{X}\right|^{2}\operatorname{polylog}\left(t_{\mathsf{mix}},\frac{1}{\xi},\frac{1}{\delta}\right),$
where $t_{\mathsf{mix}}$ is the mixing time of $P$, is sufficient in order to obtain an estimate $\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)$ of the average-mixing time $t^{\sharp}_{\mathsf{mix}}(\xi)$ which satisfies
$\begin{split}\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)\in\left[t_{\mathsf{mix}}^{\sharp}(3\xi/2),t_{\mathsf{mix}}^{\sharp}(\xi/2)\right],\end{split}$
with probability at least $1-\delta$. In other words, one can demonstrably estimate with high probability the average-mixing time linearly in the number of parameters of the transition operator, answering a question posed by Luca Zanetti and John Sylvester (Zanetti and John, 2023). This discovery stands in stark contrast with the worst-case mixing time estimation problem, or the relaxation time estimation problem, for which the sample complexity necessarily depends inversely on the minimum stationary probability (Hsu et al., 2019; Wolfer and Kontorovich, 2024) (refer to Section 4.3).
In fact, contrary to the worst-case mixing time estimation problem, we show that it is possible to extend the problem and upper bounds on the sufficient trajectory length to denumerable spaces under additional suitable conditions. Namely, when $\mathcal{X}\cong\mathbb{N}$, we show (refer to Theorem 3.5.2) that in the uniformly ergodic setting, a single path of length
$n\geq\frac{t_{\mathsf{mix}}}{\xi^{2}}\mathcal{J}_{\infty,\xi/2}\operatorname{polylog}\left(t_{\mathsf{mix}},\frac{1}{\xi},\frac{1}{\delta}\right),$
is sufficient, where for any $p\in[1,\infty]$, $\mathcal{J}_{p,\xi}$ is an entropic term defined as
$\mathcal{J}_{p,\xi}\doteq\sup\left\{\left\|Q^{(s)}\right\|_{(1-1/p)/2}^{1-1/p}\colon s\in\mathbb{N},s\leq t^{\sharp}_{\mathsf{mix}}(\xi)\right\},$
and where for $x,x^{\prime}\in\mathcal{X}$, $Q^{(s)}(x,x^{\prime})=\pi(x)P^{s}(x,x^{\prime})$. We show that one can even analyze and obtain a bound on the sufficient trajectory length in the non uniformly ergodic setting (refer to Theorem 3.5.1).
@@ 1.3  Notation and setting
$\ell_{p}\doteq\left\{v\in\mathbb{R}^{\mathcal{X}}\colon\left\|v\right\|_{p}<\infty\right\}.$
The set of all probability distributions over $\mathcal{X}$, written $\mathcal{P}(\mathcal{X})$, corresponds to non-negative functions summing to unity. The dynamics of a time-discrete, time-homogeneous Markov chain
$X_{1},X_{2},X_{3},\dots$
on $\mathcal{X}$ are governed by a transition operator represented by an infinite stochastic matrix $P$
$P(x,x^{\prime})=\mathbb{P}\left(X_{t+1}=x^{\prime}|X_{t}=x\right),\forall x,x^{\prime}\in\mathcal{X}.$
We write $\mathcal{W}(\mathcal{X})$ for the set of all infinite stochastic matrices over $\mathcal{X}$—and we also identify $\mathcal{W}(\mathcal{X})$ with Markov chains whose transition operator belong to $\mathcal{W}(\mathcal{X})$. For a finite trajectory of length $n\in\mathbb{N}$, started with an initial distribution $\mu\in\mathcal{P}(\mathcal{X})$ and evolving according to $P\in\mathcal{W}(\mathcal{X})$, the probability of observing $(x_{1},x_{2},\dots,x_{n})\in\mathcal{X}^{n}$ can be factored as
@@
Matrix multiplication retains its meaning in the countable setting. Namely, for $\mu\in\mathcal{P}(\mathcal{X})$ and $P\in\mathcal{W}(\mathcal{X})$, we write,
@@
Extensions of notation to $Pf^{\intercal}$ with $f\in\ell_{1}$, or for $P^{t}$, are similarly defined. We say that $P$ is ergodic when $P$ is irreducible, aperiodic and positive-recurrent (Levin et al., 2009). In that case, there exists a unique stationary distribution $\pi$ verifying $\pi P=\pi$, and we can define the distance to stationarity as
@@
We write $\mathcal{W}^{\star}(\mathcal{X})$ the set of uniformly ergodic Markov chains—identified with their transition operators. We recall that when $P\in\mathcal{W}^{\star}(\mathcal{X})$ the (worst-case) mixing time, defined for any $\xi\in(0,1)$ as
$\begin{split}t_{\mathsf{mix}}(\xi)\doteq\operatorname*{arg\,min}_{t\in\mathbb{N}}\left\{d(t)\leq\xi\right\},\end{split}$
is finite. It is furthermore customary to define
@@
Elements related to the notion of average-mixing time are deferred to the Section 2 for clarity.
1.4  Outline
Section 2 introduces the average-mixing time and illustrates its significance. Section 3 is dedicated to the problem of inferring the $\beta$-mixing coefficients and average-mixing time from a single trajectory of observations. In Section 3.1, we estimate the $\beta$-mixing coefficients of the chain individually, first under general ergodicity condition and then under uniform ergodicity. In Section 3.2 we deduce upper bounds for the problem of estimating the average-mixing time from a single trajectory. In Section 4, we demonstrate that our contribution holds implications at various scales of state spaces and provide more explicit rates under additional structural assumptions on the underlying chain.
Technical proofs, extensions to non-stationary chains and a discussion on the connection with $\beta$-mixing have been deferred to the supplementary material for readability.